OpenStudy (anonymous):

How would I get the standard deviation from this set of numbers? 17, 13, 13, 22, 11, 20

OpenStudy (campbell_st):

here is a site that requires you to input the data to get the summary statistics. hope it helps http://easycalculation.com/statistics/standard-deviation.php

OpenStudy (campbell_st):

otherwise its population standard deviation \[\sigma = \sqrt{\frac{\sum (x - \mu)}{N}}\] u is the mean X is the data score, N = number of scores or sample \[s = \sqrt{\frac{\sum(x - mean)}{n - 1}}\]

OpenStudy (texaschic101):

http://www.mathsisfun.com/data/standard-deviation.html

OpenStudy (anonymous):

thank you so much! It really did!